# Randomization Tests for Conditional Group Symmetry

### v1

@@ B.5 Power lower bounds for adaptive statistics
Proposition B.8.
Let $T$ be the supremum kernel distance using $\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}$ and exact randomization in the CRT. Assume that for each $\ell\in\{1,\dotsc,L\}$, $\Delta_{n,\ell}\geq 2\nu_{\ell}\sqrt{2/n}$. Then for every $\eta>0$, for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-Le^{-\eta}$ over $(Z_{1:n},{Z}^{(0)}_{1:n})$,
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})\geq F_{B,p_{\eta}(\Delta_{n}^{\vee})}(\lfloor\alpha(B+1)-1\rfloor)\;,$
$p_{\eta}(\Delta_{n}^{\vee})=\min\left\{1,L\exp\left(-\frac{3}{2}\Delta_{n}^{\vee}+R_{n}^{\vee}(\eta)\right)\right\}\;.$
@@
Using this, a lower bound for $\Delta_{n}^{\vee}$ that guarantees power $\beta$ with high probability, analogous to the one in Corollary 6.2, can be obtained.
Proof of Proposition B.8.
Recall that using the V-statistic to estimate the MMD,
$\widehat{\mathrm{SK}}(Z_{1:n},Z_{1:n}^{\prime})$ $=\sup_{\ell\in\{1,\dotsc,L\}}\left\{\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right\}$
$=\sup_{\ell\in\{1,\dotsc,L\}}\left\{\left|\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right|\right\}^{2}\;.$
Therefore, by a Fréchet inequality and Lemma B.5, and denoting
$\mathcal{A}_{1:n}=(S_{1:n},Z_{1:n},Z^{(0)}_{1:n})$,
$\mathbb{P}\left(\widehat{\mathrm{SK}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$=1-\mathbb{P}\left(\widehat{\mathrm{SK}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\leq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
@@
$\leq\min\left\{1,\sum_{\ell=1}^{L}\mathbb{P}\left(\left|\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\right|\leq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}\mid\mathcal{A}_{1:n}\right)\right\}$
$=\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\frac{\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})}{3a_{\ell}+\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}+\frac{9}{2}a_{\ell}+\frac{3}{2}\bar{H}_{\ell}\right)\right\}\;.$
Using $a^{\vee}$ and $\bar{H}^{\vee}$ as defined above,
@@
$\mathbb{P}\left(\widehat{\mathrm{SK}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})\mid S_{1:n},Z_{1:n},{Z}_{1:n}^{(0)}\right)$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}+\frac{9}{2}a^{\vee}+\frac{3}{2}\bar{H}^{\vee}\right)\right\}\;.$
Thus, for any $\eta>0$, with probability at least $1-Le^{-\eta}$,
@@
$\mathbb{P}\left(\widehat{\mathrm{SK}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{SK}}(Z_{1:n},{Z}_{1:n}^{(0)})\mid S_{1:n}\right)$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\Delta_{n}^{\vee}+R_{n}^{\vee}(\eta)\right)\right\}\;,$
where $\Delta_{n}^{\vee}=\sup_{\ell\in[L]}\Delta_{n,\ell}$, and $R_{n}(\eta)$ is as defined above. The rest of the proof is analogous to the proof of Theorem 6.1. ∎
Appendix C Examples of exact conditional randomization
@@
Here, we describe two possible assumptions for $P_{\Gamma|[X]}$ that each allow for sampling from the exact conditional distribution.
@@
$P_{\Gamma|[X]}=P_{\Gamma}$ for all $M(X)$.
Assumption 1 says that the distribution over the group action is the same across orbits. Generating a new sample $\mathsf{\Gamma}_{1},\dotsc,\mathsf{\Gamma}_{n}$ can then be done by simply permuting $\tilde{\tau}_{1},\dotsc,\tilde{\tau}_{n}$. The resulting randomization procedure using Property E3 is outlined in Algorithm 4. We note that this assumption is trivially true if $\mathbf{G}$ acts transitively on $\mathbf{X}$.
@@
1:sample $\tilde{\tau}_{1},\dotsc,\tilde{\tau}_{n}$ with $\tilde{\tau}_{i}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X_{i})$
2:sample permutation $\tilde{\tau}_{\pi(1)},\dotsc,\tilde{\tau}_{\pi(n)}$
3:compute $\left((\tilde{\tau}_{\pi(1)}\rho(X_{1}),\tilde{\tau}_{\pi(1)}\tilde{Y}_{1}),\dotsc,(\tilde{\tau}_{\pi(n)}\rho(X_{n}),\tilde{\tau}_{\pi(n)}\tilde{Y}_{n})\right)$
@@
Suppose there is a group $\mathbf{H}$ acting transitively on $\mathbf{M}$, and let $\phi\colon\mathbf{H}\to\mathbf{G}$ be a group homomorphism. Then $P_{\Gamma|[X]}$ is $\mathbf{H}/\mathbf{G}$-equivariant in the sense that for all $h\in\mathbf{H}$,
$P_{\Gamma|[X]}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid hM(X))=P_{\Gamma|[X]}(\phi(h)^{-1}{\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid M(X))\;.$
Assumption 2 says that the conditional distribution over the group action is itself equivariant with respect to the action of $\mathbf{H}$ on $\mathbf{M}$. Assuming that $\phi$ is known, new samples can be generated by permuting $\tilde{\tau}_{1},\dotsc,\tilde{\tau}_{n}$ and applying a corresponding transform under $\phi$. The corresponding randomization procedure using Property E3 is outlined in Algorithm 5.
@@
1:compute $\tilde{\tau}_{1},\dotsc,\tilde{\tau}_{n}$ with $\tilde{\tau}_{i}\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid X_{i})$
2:sample permutation $\tilde{\tau}_{\pi(1)},\dotsc,\tilde{\tau}_{\pi(n)}$
@@
4:compute $\left((\phi(\tilde{h}_{\pi(i)})^{-1}\tilde{\tau}_{\pi(i)}\rho(X_{i}),\phi(\tilde{h}_{\pi(i)})^{-1}\tilde{\tau}_{\pi(i)}\tilde{Y}_{i})\right)_{i=1}^{n}$
Appendix D Additional experiments and experimental details
D.1 Test for invariance p-values
Figure 9 shows histograms of the $p$-values for the tests for invariance described in Example 1 from Section 2.2. In particular, the distributions of $p$-values for the randomization tests are approximately uniform.
Figure 9: Histograms showing the distribution of $1000$ $p$-values for the tests described in Example 1 when $\theta$ is sampled uniformly from $[0,p)$. Labels are the same as in Figure 2.
D.2 Additional results for the experiment in Section 7.2
@@ D.1 Test for invariance p-values
For the first experiment discussed in Section 7.2, Figure 10 shows the estimated rejection rates for the tests that use independent comparison sets. For the second experiment where non-$\mathrm{SO}(d)$-equivariance is isolated to the first two dimensions, Figure 11 shows the estimated rejection rates for tests that use independent comparison sets. In both experiments, results for the tests that use independent comparison sets are similar to that of tests that reuse a set.
@@
Figure 10: Rejection rates for tests for $\mathrm{SO}(d)$-equivariance that use independent comparison sets in the changing covariance experiment for varying values of the $Y|X$ covariance matrix off-diagonal values $p$ and sample size $n$. Dotted lines show the corresponding rates for tests that reuse a comparison set.
Figure 11: Rejection rates for tests that use independent comparison sets in the experiment where the non-SO(d)-equivariance comes from the covariance $p$ between the first two dimensions, for varying values of the sample size $n$. Dotted lines show the corresponding rates for tests that reuse a comparison set.
D.3 Effect of the number of randomizations
In the same setting described in Section 7.2 that takes the conditional distribution to be non-equivariant in all dimensions, we investigate the effect of the number of randomizations $B$ used to estimate the $p$-value. Figure 12 shows the estimated rejection rates for the equivariant and a non-equivariant setting as $B$ ranges between $50$ and $1000$. In the equivariant setting, the rejection rates remain well-under the expected rate of $0.05$. In the non-equivariant setting, the power remains relatively consistent after $B=100$ or so.
Figure 12: Rejection rates for the changing covariance experiment for number of randomizations $B$ ranging from $50$ to $1000$. For each set of tests, the conditional distribution is $\mathrm{SO}(3)$-equivariant for the plots in the first row and not equivariant for the plots in the second.
D.4 Sampling from approximation versus true conditional distribution
Consider a setting where $X$ is generated in the following way. Let $Z\sim\chi_{1}$ be a chi random variable with $1$ degree of freedom, and take $X\coloneq Z\Gamma\mathbf{e}_{1}$, where $\Gamma|Z\sim P_{\tilde{\tau}|X}({\,\vbox{\hbox{\tiny$\bullet$}}\,}|Z1_{d}+\varepsilon)$ is a $d$-dimensional rotation matrix, with $1_{d}$ being the vector of ones of length $d$ and $\varepsilon\sim\mathsf{N}(0,\mathbf{I}_{d})$. In other words, the conditional distribution $P_{\Gamma|[X]}$ is known and can be sampled from exactly. Let $Y|X$ have the conditional distribution described in Section 7.2, with $p$ being the covariance between any pair of dimensions of $Y$. Recall that the conditional distribution of $Y|X$ is $\mathrm{SO}(d)$-equivariant if and only if $p=0$.
For $d\in\{3,4,5\}$, we compare the performance of tests that sample from an approximation of $P_{\Gamma|[X]}$ as per Algorithm 2 with tests that sample from $P_{\Gamma|[X]}$ exactly. The test results are shown in Figure 13. In this setting, we see that kernel methods well-approximate sampling from the true conditional. Histograms showing the distribution of $p$-values for these tests are provided in Figures 14 and 15. In the equivariant setting $(p=0)$, we see that the distributions of $p$-values for the conditional randomization tests are approximately uniform for both tests that use approximate sampling and tests that use exact sampling.

### v2

@@ C.5 Power lower bounds for adaptive statistics
Proposition C.10.
Let $T$ be the supremum kernel distance using $\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}$ and exact randomization in the CRT. For any $\eta>0$, if
$\Delta^{\lor}_{n}>2\kappa_{\lor}\left(2\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\;,$ (27)
then for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $\max\{0,1-Le^{-\eta}\}$ over $(Z_{1:n},{Z}^{(0)}_{1:n})$,
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})\geq F_{B,p_{\eta}(\Delta^{\lor}_{n})}(\lfloor\alpha(B+1)-1\rfloor)\;,$
where $F_{B,p_{\eta}(\Delta^{\lor}_{n})}$ is the distribution function for the binomial distribution of $B$ trials with success probability
$p_{\eta}(\Delta^{\lor}_{n})=\min\left\{1,L\exp\left(-\frac{3}{2}\Delta^{\lor}_{n}+C_{n}^{\vee}(\eta)\right)\right\}\;.$
@@
Using this, a lower bound for $\Delta_{n}^{\vee}$ that guarantees power $1-\beta$ with high probability, analogous to the one in Corollary C.8, can be obtained.
Proof of Proposition C.10.
Let
$T(Z_{1:n},Z_{1:n}^{\prime})$ $\coloneqq\widehat{\mathrm{SK}}(Z_{1:n},Z_{1:n}^{\prime})$
$=\sup_{\ell\in[L]}\widehat{\mathrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})$
$=\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right|^{2}\;,$
and let $\widehat{T}\coloneqq T(Z_{1:n},Z_{1:n}^{(0)})$ denote the observed test statistic. Denote $\mathcal{A}_{1:n}\coloneqq(S_{1:n},Z_{1:n},Z_{1:n}^{(0)})$. Observe that
[…]
@@
$=\mathbb{P}\left(\left.\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|\geq\sqrt{\widehat{T}}\;\right|\;\mathcal{A}_{1:n}\right)$
@@
$=\mathbb{P}\left(\left.\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n}^{\lor}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n}\right)$
@@
$\leq\min\left\{1,\mathbb{P}\left(\left.\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n}^{\lor}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n},\sqrt{\widehat{T}}>\xi_{n}^{\lor}\right)\right$.
$\qquad\qquad+\mathbbm{1}\left\{\sqrt{\widehat{T}}\leq\xi_{n}^{\lor}\right\}\Bigg\}\;,$
where the last inequality follows from the same argument as that in Theorem 5.1. We first derive a high-probability upper bound for the first term. Applying a Fréchet inequality (first inequality), Lemma C.6 (third inequality), and the same factoring trick as in (22) (fourth inequality),
$p^{*}$ $\coloneqq\mathbb{P}\left(\left.\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n}^{\lor}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n},\sqrt{\widehat{T}}>\xi_{n}^{\lor}\right)$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\mathbb{P}\left(\left.\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n}^{\lor}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n},\sqrt{\widehat{T}}>\xi_{n}^{\lor}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\mathbb{P}\left(\left.\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n,\ell}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n},\sqrt{\widehat{T}}>\xi_{n}^{\lor}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\frac{\left(\sqrt{\widehat{T}}-\bar{H}_{n,\ell}\right)^{2}}{3a_{\ell}+\sqrt{\widehat{T}}-\xi_{n}^{\lor}}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\sqrt{\widehat{T}}+\frac{3}{2}\xi_{n}^{\lor}+\frac{9}{2}a_{\ell}\right)\right\}$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\sqrt{\widehat{T}}+\frac{3}{2}\xi_{n}^{\lor}+\frac{9}{2}a^{\lor}\right)\right\}\;.$
Applying a union bound with (23) and the bounds on $a^{\lor}$ (25) and $\xi_{n}^{\lor}$ (26), for any $\eta>0$, with probability at least $\max\{0,1-Le^{-\eta}\}$ over $Z_{1:n}$ and $Z_{1:n}^{(0)}$,
$\exp\left(-\frac{3}{2}\sqrt{\widehat{T}}+\frac{3}{2}\xi_{n}^{\lor}+\frac{9}{2}a^{\lor}\right)$
$\quad\leq\exp\left\{-\frac{3}{2}\left(\Delta^{\lor}_{n}-2\kappa_{\lor}\left(\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\right)\right$.
$\quad\qquad\qquad\left.+\frac{3}{2}\left(2\kappa_{\lor}\sqrt{\frac{2}{n}}\right)+\frac{9}{2}\kappa_{\lor}^{2}\left(\frac{2}{n}+16\sqrt{\frac{2}{n^{3}}}\right)\right\}$
$\quad=\exp\left(-\frac{3}{2}\Delta^{\lor}_{n}+C_{n}^{\lor}(\eta)\right)\;.$
Hence,
$p^{*}\leq\min\left\{1,L\exp\left(-\frac{3}{2}\Delta^{\lor}_{n}+C_{n}^{\lor}(\eta)\right)\right\}=:p_{\eta}(\Delta^{\lor}_{n})\;.$
For the indicator term in the upper bound of p(T^)p($\widehat${T}), by (23) and (26), with the same probability,
$\mathbbm{1}\left\{\sqrt{\widehat{T}}\leq\xi_{n}^{\lor}\right\}$ $=\mathbbm{1}\left\{\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{(0)})\right|\leq\xi_{n}^{\lor}\right\}$
$\leq\mathbbm{1}\left\{\Delta^{\lor}_{n}-2\kappa_{\ell}\left(\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\leq\xi_{n}^{\lor}\right\}$
$\leq\mathbbm{1}\left\{\Delta^{\lor}_{n}\leq 2\kappa_{\ell}\left(2\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\right\}\;,$
which is zero under assumption (27). Thus, $p(\widehat{T})\leq p_{\eta}(\Delta^{\lor}_{n})$ with probability at least $\max\{0,1-Le^{-\eta}\}$.
The rest of the proof is analogous to the second part of Theorem 5.1.
Appendix D Further considerations for randomization tests for conditional symmetry
D.1 Randomization using representative inversions with non-free group actions
As demonstrated in Example 2, a representative inversion $\psi^{\circ}$} may be (non-uniquely) defined even when the group action is not free in some cases. In these cases, a test for conditional symmetry that uses the deterministic representative inversion is still a valid level-$\alpha$ test under the null. This follows from the observation that for a random action $\Psi|X\stackrel{{\scriptstyle\text{\rm\tiny d}}}{{=}}{\psi^{\circ}}(X)H|X$, where $H$ is a random element of the stabilizer subgroup $\mathbf{G}_{X}$ of $X$, assuming that $Y|X$ is $\mathbf{G}$-equivariant,
$P_{Y|X}(\Psi^{-1}Y\in B\mid x)$ $=P_{Y|X}(({\psi^{\circ}}(x)H)^{-1}Y\in B\mid x)$
$=P_{Y|X}({\psi^{\circ}}(x)^{-1}Y\in B\mid H^{-1}x)=P_{Y|X}({\psi^{\circ}}(x)^{-1}Y\in B\mid x)\;.$
That is, the conditional distributions of $\Psi^{-1}Y|X$ and ${\psi^{\circ}}(X)^{-1}Y|X$ are the same. A test that uses the representative inversion may be more computationally efficient than one that samples from the inversion kernel. However, tests will generally have different properties under alternatives. We explore such differences empirically in Section E.4.
D.2 Exact conditional randomization
@@
Here, we describe two possible assumptions for $P_{\Gamma|M}$ that each allow for sampling from the exact conditional distribution.
@@
$P_{\Gamma|M}=P_{\Gamma}$ for all $M(X)$.
Assumption 1 says that the distribution over the group action is the same across orbits. Generating a new sample $\mathsf{\Gamma}_{1},\dotsc,\mathsf{\Gamma}_{n}$ can then be done by simply permuting $\Psi_{1},\dotsc,\Psi_{n}$. The resulting randomization procedure using Property E3 is outlined in Algorithm 4. We note that this assumption is trivially true if $\mathbf{G}$ acts transitively on $\mathbf{X}$.
@@
1:sample $\Psi_{1},\dotsc,\Psi_{n}$ with $\Psi_{i}|X_{i}\sim P_{\Psi|X}$
2:sample permutation $\Psi_{\pi(1)},\dotsc,\Psi_{\pi(n)}$
3:compute $\left((\Psi_{\pi(1)}\rho(X_{1}),\Psi_{\pi(1)}\tilde{Y}_{1}),\dotsc,(\Psi_{\pi(n)}\rho(X_{n}),\Psi_{\pi(n)}\tilde{Y}_{n})\right)$
@@
Suppose there is a group $\mathbf{H}$ acting transitively on $\mathbf{M}$, and let $\phi\colon\mathbf{H}\to\mathbf{G}$ be a group homomorphism. Then $P_{\Gamma|M}$ is $\mathbf{H}/\mathbf{G}$-equivariant in the sense that for all $h\in\mathbf{H}$,
$P_{\Gamma|M}({\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid hM(X))=P_{\Gamma|M}(\phi(h)^{-1}{\,\vbox{\hbox{\tiny$\bullet$}}\,}\mid M(X))\;.$
Assumption 2 says that the conditional distribution over the group action is itself equivariant with respect to the action of $\mathbf{H}$ on $\mathbf{M}$. Assuming that $\phi$ is known, new samples can be generated by permuting $\Psi_{1},\dotsc,\Psi_{n}$ and applying a corresponding transform under $\phi$. The corresponding randomization procedure using Property E3 is outlined in Algorithm 5.
@@
1:compute $\Psi_{1},\dotsc,\Psi_{n}$ with $\Psi_{i}|X_{i}\sim P_{\Psi|X}$
2:sample permutation $\Psi_{\pi(1)},\dotsc,\Psi_{\pi(n)}$
@@
4:compute $\left((\phi(\tilde{h}_{\pi(i)})^{-1}\Psi_{\pi(i)}\rho(X_{i}),\phi(\tilde{h}_{\pi(i)})^{-1}\Psi_{\pi(i)}\tilde{Y}_{i})\right)_{i=1}^{n}$
Appendix E Additional experiments and experimental details
E.1 Using independent comparison sets
@@
For the experiment discussed in Section 6.2.1, Figure 8 shows the estimated rejection rates for the tests that use independent comparison sets. We observe that the results for the tests that use independent comparison sets are similar to those of tests that reuse a single comparison set.
@@
Figure 8: Rejection rates for tests for $\mathrm{SO}(d)$-equivariance that use independent comparison sets in the changing covariance experiment for varying values of the $Y|X$ covariance matrix off-diagonal values $c$ and sample size $n$. Dotted lines show the corresponding rates for tests that reuse a comparison set. (Note the different vertical scale for $c=0$ versus $c>0$.)
E.2 Approximation via kernel density estimation
Figure 9: Rejection rates for $\mathrm{SO}(d)$-equivariance RN-FS and RN-SK tests for varying values of the $Y|X$ covariance matrix off-diagonal values $c$ and sample size $n$. The number suffix in the test labels indicates the constant scaling applied to the bandwidth on the maximal invariant kernel used for conditional randomization, where the test with a scaling of $1.0$ is the one used in other experiments. (Note the different vertical scale for $c=0$ versus $c>0$.)
For the same setting as in Section 6.2.1, Figure 9 shows how approximating the conditional randomization via kernel methods can change the size and power of the FUSE and SK tests. Here, we apply different constant scalings to the bandwidth on the kernel used in the approximate conditional randomization method described in Section 4.1, where the default bandwidth follows Silverman’s rule of thumb for multivariate kernels (Silverman, 1986, Ch. 4). The rule says to take the bandwidth matrix $\mathbf{H}$ as a diagonal matrix with the diagonal elements being
$\left(\frac{4}{d+2}\right)^{\frac{2}{d+4}}n^{\frac{-2}{d+4}}(s\hat{\sigma}_{j})^{2}\;,$
where $d$ is the dimension of $M$, $\hat{\sigma}^{2}$ is the sample variance in the $j$-th dimension, and $s=1$ is the default scaling. In our setting, we observe that with less smoothing ($s\in\{0.1,1,2\}$), the test is conservative with a rejection rate less than $\alpha=0.05$, while over-smoothing ($s\in\{5,10\}$) can increase power but also inflate Type I error.
E.3 Risks in random inference
Figure 10: Histograms for the estimated resampling risk (RR) and risk of overestimated significance (ROS) for RN-FS in testing for $\mathrm{SO}(d)$-equivariance with $B=100$ and $n=250$ across 100 simulations.
For the FUSE test for $\mathrm{SO}(d)$-equivariance in the setting described in Section 6.2.1, we estimate its resampling risk (RR) and risk of overestimated significance (ROS) (Stoepker and Castro, 2024). These quantities measure risks in inference that arise purely due to the Monte Carlo estimation of the $p$-value. Given a fixed dataset $z_{1:n}$, let $p^{*}(z_{1:n})$ denote the “true” $p$-value (the value obtained from $p_{B}$ in (11) by taking the number of randomizations $B\to\infty$), and let $\phi^{*}(z_{1:n})\in\{0,1\}$ denote the result of the test based on $p^{*}(z_{1:n})$. The RR measures the probability of obtaining a test decision differing from the one based on $p^{*}(z_{1:n})$ as a consequence of the Monte Carlo estimation:
$\mathrm{RR}(z_{1:n})=\mathbb{P}\left\{\phi_{B}(z_{1:n})\neq\phi^{*}(z_{1:n})\right\}\;.$
The ROS measures the probability of obtaining a $p$-value smaller than $p^{*}(z_{1:n})$:
$\mathrm{ROS}(z_{1:n})=\mathbb{P}\left\{p_{B}(z_{1:n})<p^{*}(z_{1:n})\right\}\;.$
In this experiment, we estimate $p^{*}(z_{1:n})$ by running the test with $B^{*}=10,000$ and using that estimate to compute the above theoretical probabilities for a test with $B=100$. Figure 10 shows histograms of the estimated RR and ROS across $N=100$ independent datasets. Overall, RR remains low across datasets, indicating that random inference does not greatly affect the result of the test. On the other hand, the ROS may be nontrivial at times, suggesting that the $p$-values obtained from the test may not be interpretable in terms of significance without using a larger $B$.
E.4 Randomization using representative inversions
Consider the same experimental setup as in Section 6.2.1 but with tests that directly use the representative inversion defined in (4) as opposed to sampling actions from the inversion kernel $P_{\Psi|X}$. The results are shown in Figure 11. We see that these tests retain their approximate level-$\alpha$ size when the true conditional distribution is equivariant $(c=0)$. We also observe that these tests have slightly different power properties under the alternative compared to those that randomize using the inversion kernel. Figure 12 show the corresponding $p$-value histograms for these tests.